Determine the solutions to the equation cosx/sin2x =5/7 for 0 less than equal to x less than equal to 2pi accurate to two decimal places.
Answer by Boreal(15235)

sin 2x=2 sin x * cos x
cos x/sin 2x=cos x/2 sin x* cos x=1/(2 sin x)
1/(2 sin x)=5/7
10 sin x=7
sin x=(0.7)
arc sin of both sides, and x=44.4 degrees
cos 44.4=0.714
sin 88.8=1 (essentially)
0.714/1=5/7